1.) Find the slope of the following line.
4y=x+4
The slope is?


2.) Give the slope of each line and determine whether the two are parrallel, perpindicular, or neither.
5x+2y=2
3x+7y=0
The slope of the first line is?
The slope of the second line is?
Are they parrallel, perpindicular, are neither?
3.) Find the slope-intercept of the line that has the given characteristics. Slope (0) and y-intercept (0.8).
The equation is y=?

4.) Use the slope and y-intercept if possible, to graph the line whose equation is given.
y=4x+3


5.) Graph the equation using the slope and the y-intercept.
x+5y=20

Answer by mananth(16946)

1.) Find the slope of the following line.
4y=x+4
y=(1/4)x+1
compare with y= mx+b
m= slope = 1/4
The slope is?


2.) Give the slope of each line and determine whether the two are parrallel, perpindicular, or neither.
5x+2y=2
2y=-5x+2
/2
y=(-5/2)x+1
slope = (-5/2)
3x+7y=0
7y= -3x
y=(-3/7)x
slope = (-3/7)
Neither parallel nor perpendicular

3.) Find the slope-intercept of the line that has the given characteristics. Slope (0) and y-intercept (0.8).
The equation is y=? y=8

4.) Use the slope and y-intercept if possible, to graph the line whose equation is given.
y=4x+3




5.) Graph the equation using the slope and the y-intercept.
x+5y=20